OpenStudy (anonymous):

how do you get the square root of a binomial out of this problem x^2+20x+100

OpenStudy (compassionate):

You want to factor each terms using FOIL

OpenStudy (anonymous):

our teacher didnt teach us how.

OpenStudy (compassionate):

Okay, so first thing; do you know how to factor?

OpenStudy (anonymous):

somewhat

OpenStudy (compassionate):

Say, if I gave you (a + b)(2 + 2) Could you factor it?

OpenStudy (anonymous):

i doubt it i wouldnt now how to factor that. not to good at math.

OpenStudy (compassionate):

Fine, you use the foil method. FOIL = First, outer, Inner, Last. x^2+20x+100 ANS: (x + 10 )(x + 10 ) FOIL x^2 + 10x+10x + 100 x^2 + 20x + 100, seems to be okay. Next time I suggest not being so rude when someone is trying to help you.

OpenStudy (compassionate):

Oh, pardon me, I misread your comment! You weren't rude. My apologies.

OpenStudy (anonymous):

its cool, i still dont understand ill give you an example of a problem in my book so you will know that your doing it right. x^2+10x-3=0 then it shows x^2+10x=3. then x^2+10x+5^2=3+25 then (x+5)=square root of 28 then x=-5 square root of 28 then x=5 plus or minus 2 square root of 7

OpenStudy (compassionate):

Okay, and what exactly are you having trouble with?

OpenStudy (compassionate):

The equation part? PEMDAS part? Factoring? Expanding?

OpenStudy (anonymous):

all of it my teacher, sucks.

OpenStudy (anonymous):

i know to write the left side as x^2+bx which is x^2+20x=100

OpenStudy (compassionate):

Give me a problem and I'll walk you through.

OpenStudy (anonymous):

the first problem i gave u please

OpenStudy (compassionate):

I already solved it... You want me to do step by step on how I solved it?

OpenStudy (anonymous):

yes please

OpenStudy (compassionate):

I found this: http://www.themathpage.com/alg/quadratic-trinomial.htm It's very helpful.

OpenStudy (anonymous):

thank you

OpenStudy (compassionate):

Go down to the quadratic trinomial part

OpenStudy (compassionate):

I mean Multiplying Binomials, then read

OpenStudy (anonymous):

would the answer be x=10-10 square root of 20?

OpenStudy (compassionate):

For: x^2+20x+100?

OpenStudy (anonymous):

yes

OpenStudy (compassionate):

We pull put them in parenthesis. ( )( ) <--- First we factor x^2, do you know how we can do that?

OpenStudy (anonymous):

from the examples in my book none of the answer are in parenthesis thats where you confusing me

OpenStudy (compassionate):

Here, read this and then come back to me with questions or concerns: http://www.themathpage.com/alg/factoring-trinomials.htm

OpenStudy (compassionate):

Not in parenthesis? Give me an example.

OpenStudy (anonymous):

its to complicated to type out i really appreciate the effort your giving to try to help me but i cant explain what were doing therefore theres not a whole lot you can do

OpenStudy (compassionate):

If you give me a problem I can tell you what is happening

OpenStudy (anonymous):

its hard to type it out

OpenStudy (compassionate):

Well, can you try to explain it?

OpenStudy (anonymous):

i have tried

OpenStudy (compassionate):

Hmm... Well, then I cannot help you further. I really do want to help, but if you can't even explain it then your teacher is doing a bad job. Report this to the principle, tell him/her that your teacher isn't doing a good job.

OpenStudy (anonymous):

she has been at the school for years my mom had her and she still isnt gone

OpenStudy (compassionate):

Keep reporting.

OpenStudy (anonymous):

its been like 21 years since my mom went to school and shes still there, there is no hope

OpenStudy (compassionate):

Ohlawd get a new school?

OpenStudy (anonymous):

wish i could trust me

OpenStudy (compassionate):

Tell your teacher that she's horrible and she should feel bad.

OpenStudy (anonymous):

she doesnt even care. i swear ask her one question she freaks out says she doesnt have time to help you and to read the chapter. shes a b*

OpenStudy (compassionate):

Kill her.

OpenStudy (anonymous):

yeah right

OpenStudy (compassionate):

Lol. It was a joke

OpenStudy (anonymous):

i know lol

OpenStudy (compassionate):

Well, can I get a best answer? ;D

OpenStudy (anonymous):

u got it lol

OpenStudy (compassionate):

Thanks xD

OpenStudy (anonymous):

no problem lol